Mathematics 241 – Multivariable Calculus Quiz 7 – November 8, 2013

Your Name:

A) (20) Using the method of Lagrange Multipliers, find the maximum and minimum values of f(x, y) = 2x + 3y on the constraint curve

E = \{(x, y) | x^{2} + \frac{y^{2}}{9} = 1\}.

(If you run into trouble with the algebra, for more partial credit, sketch what the level curves of f should look like at the maximum and minimum together with the constraint curve.)